OpenStudy (anonymous):

* Advance Algebra With Finical Application * Explain how you would calculate the amount of your payment to the government or the amount of your refund. Provide the solution ----------------------

OpenStudy (anonymous):

My work ! ----------- Part I Scenario 1 Part II 49,500-1,800=44700 I will choose itemized deduction. AGI is 47700 Value of exemption is 3650 federal withholding 2985-750-2700= 6435 44050-6435=37615 Part III 837.50+[15%*(49,500-8,375.00)] =837.50+[0.15*$41125] =837.50+[6168.75] =7006.25 Part IV I